Given a random sample of 5 pints from different
production lots, we want to test whether the fat content
of a certain kind of ice cream exceeds 14%. What can
be conclude at the 0.01 level of significance about the
null hypothesis µ = 14% if the Sample has mean
X = 14.9% and the standard deviation σ = 0.42%?

Answer by stanbon(75887)

Given a random sample of 5 pints from different
production lots, we want to test whether the fat content
of a certain kind of ice cream exceeds 14%. What can
be conclude at the 0.01 level of significance about the
null hypothesis µ = 14% if the Sample has mean
X = 14.9% and the standard deviation σ = 0.42%?
---------------
Ho:: p >= 0.14
Ha: p < 0.14
----------------
x-bar = 0.149
---
z(0.149) = (0.149-0.14)/sqrt[0.42*0.58/5] = 0.0408
-----
p-value = P(z < 0.0408) = normalcdf(-100,0.0408) = 0.5163
--------
Since the p-value is greater than 1% fail to reject Ho.
------------
Cheers,
Stan H.
--------------